### v2

# Equivariant Spherical Transformer for Efficient Molecular Modeling

Abstract
SE(3)-equivariant Graph Neural Networks (GNNs) have significantly advanced molecular system modeling by employing group representations. However, their message passing processes, which rely on tensor product-based convolutions, are limited by insufficient non-linearity and incomplete group representations, thereby restricting expressiveness. To overcome these limitations, we introduce the Equivariant Spherical Transformer (EST), a novel framework that leverages a Transformer structure within the spatial domain of group representations after Fourier transform. We theoretically and empirically demonstrate that EST can encompass the function space of tensor products while achieving superior expressiveness. Furthermore, EST’s equivariant inductive bias is guaranteed through a uniform sampling strategy for the Fourier transform. Our experiments demonstrate state-of-the-art performance by EST on various molecular benchmarks, including OC20 and QM9.
1 Introduction
Graph neural networks (GNNs) have become increasingly prevalent for modeling molecular systems and approximating quantum mechanical calculations, providing crucial support for various computational chemistry tasks, including drug discovery senior2020improved and material design zitnick2020introduction . Compared to traditional calculation methods like Density Functional Theory (DFT), GNNs significantly reduce the computational cost of quantum property prediction from hours or days to fractions of a second. However, directly applying regular GNNs (e.g. GCN kipf2016semi and GIN xu2018powerful ) to 3D molecular conformations often yields poor results due to their neglect of inherent physical constraints. To overcome these limitations, equivariant GNNs have emerged as a more promising avenue, effectively capturing intricate atomic interactions and addressing multiple crucial challenges.
In molecular systems, SE(3)-invariance and SE(3)-equivariance are fundamental constraints. For instance, molecular energies remain constant under rotation, while atomic forces rotate in concert with the molecule equivariantly. Early invariant models, such as SchNet schutt2018schnet and HIP-NN lubbers2018hierarchical , rely on interatomic distances in their message-passing blocks, consequently limiting their ability to capture interactions involving triplets or quadruplets of atoms miller2020relevance . Directional GNNs gasteiger2020directional ; gasteiger2021gemnet address this problem by explicitly incorporating bond angles and dihedral angles. Despite performance improvements, their expressiveness remained constrained as they primarily extracted features from invariant and handcrafted features. To capture deeper features, SE(3)-equivariant GNNs employ group representations as node embeddings and construct steerable message-passing blocks thomas2018tensor ; brandstetter2021geometric ; liao2023equiformer , where tensor product are applied to capture equivariant interactions between group embeddings. These GNNs achieve improved performance using only raw molecular structures, without providing handcrafted features. Nevertheless, tensor product operations inherently suffer from limited non-linearity thomas2018tensor , and their expressiveness is bounded by the degree of group representations dym2021on ; joshi2023expressive . An alternative equivariant approach processes features in the spatial domain after Fourier transforming from group representations, where embeddings are defined by square-integrable spherical functions (see Figure 1(a)). Prior works zitnick2022spherical ; passaro2023reducing ; liao2024equiformerv typically apply simple point-wise neural networks to model signals on these spherical functions, operating independently on each orientation (see Figure 1(b)). Introducing nonlinearity across different orientations within these spherical functions is rarely considered due to the risk of violating equivariance. Nevertheless, we show that this operation can be performed equivariantly, potentially providing additional nonlinearity and expressiveness (Further analysis is provided in Section 3.2.)
In this paper, we present the Equivariant Spherical Transformer (EST), a novel message-passing framework for accurate atomic interaction modeling. First, we use the spherical functions in the spatial domain as node and edge embeddings, and transform it to point sequences, where each point represent a unique orientation on the spherical functions. A spherical attention mechanism is then introduced to capture dependencies within these point sequences (see Figure 1(c)). Furthermore, EST incorporates a mixture of hybrid experts structure, utilizing feed-forward networks in both spatial domain and original harmonic domain to effectively balance equivariance and model capacity. The equivariant inductive bias of EST is guaranteed through our implementation of a uniform sampling strategy in spatial domain. Our experiments show that EST-based architectures achieve state-of-the-art (SOTA) performance on various benchmarks such as OC20 chanussot2021open and QM9 ramakrishnan2014quantum , exhibiting particularly stable and significant improvements for complex molecular systems.
Our contributions are: (i) We propose EST, a novel framework whose expressiveness provably subsumes traditional tensor products. (ii) By employing an uniform sampling implementation, we guarantee the equivariance of EST. (iii) We conduct a range of experiments for quantum property prediction tasks, demonstrating superior performance of EST. Moreover, our ablation studies confirming the enhanced expressiveness and equivariance of EST.
2 Related Work and Preliminaries
In this section, we discuss related work and the corresponding mathematical background relevant to equivariant GNNs. We begin by listing the notations frequently used throughout the paper. We denote the unit sphere as $\mathbb{S}^{2}$, where the coordinate of a spherical point (or orientation) $\vec{\mathbf{p}}=(\theta,\varphi)$ is represented by its polar angle $\theta$ and azimuth angle $\varphi$. The symbol $\mathbb{R}$ represents the set of real numbers, while $\mathbf{R}$ denotes a rotation matrix for 3D vectors. We use $\cdot$$\cdot$] to indicate tensor concatenation and $\tilde{\cdot}$ to represent the update of tensors, respectively.
Figure 1: Overview of operations in the spherical spatial domain. (a) Applying the Fourier transform to group steerable representations, projecting them onto the spatial domain and storing them via spherical sampling. (b) A conventional point-wise operation on the sphere, where a function $g(\cdot)$ combines features across different spherical functions at a given orientation. (c) The proposed EST operation, where $g(\cdot)$ represents a Transformer framework designed to model global dependencies and interactions among different orientations on the sphere.
Message Passing Neural Networks
Consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with nodes $v_{i}\in\mathcal{V}$ and edges $e_{ij}\in\mathcal{E}$. Each node $v_{i}$ has an embedding $\mathbf{x}_{i}$ and an attribute $\mathbf{z}_{i}$, and each edge $e_{ij}$ has an attribute $\mathbf{a}_{ij}$. The Message Passing Neural Network (MPNN) schutt2018schnet , a specific type of GNN, updates node embeddings through a message block $\mathbf{M}(\cdot)$ and an update block $\mathbf{U}(\cdot)$ via the following steps:
$\mathbf{m}_{ij}=\mathbf{M}(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{z}_{i},\mathbf{z}_{j},\mathbf{a}_{ij},\vec{\mathbf{r}}_{ij}),\quad\text{and}\quad\tilde{\mathbf{x}}_{i}=\mathbf{U}(\mathbf{x}_{i},\sum_{j\in\mathcal{N}(i)}\mathbf{m}_{ij}),$ (1)
where we use $\vec{\mathbf{r}}_{ij}$ to denote the relative position of node $i$ and node $j$ in 3D space, and the neighborhood $\mathcal{N}(i)$ is typically defined by a cutoff radius: $\mathcal{N}(i)=\{j\mid\|\vec{\mathbf{r}}_{ij}\|\leq r_{cut}\}$. The node attribute $\mathbf{z}_{i}$ contains atomic information such as the atomic type, and the edge attribute $\mathbf{a}_{ij}$ contains atomic pair information such as distance and bond type. By stacking multiple message and update blocks, the final node embeddings can be used to model atomic interactions or represent molecular properties.
Equivariance and Invariance
Given a group $\mathbf{G}$ and a transformation parameter $g\in\mathbf{G}$, a function $\phi:\mathcal{X}\to\mathcal{Y}$ is said to be equivariant to $g$ if it satisfies:
$\phi(T(g)[x])=T^{\prime}(g)[\phi(x)],$ (2)
where $T^{\prime}(g):\mathcal{Y}\to\mathcal{Y}$ and $T(g):\mathcal{X}\to\mathcal{X}$ denote the corresponding transformations on $\mathcal{Y}$ and $\mathcal{X}$, respectively. Invariance is a special case of equivariance where $T^{\prime}(g)$ is the identity transformation. In this paper, we primarily focus on SO(3) equivariance, i.e., equivariance under 3D rotations, as it is closely related to the interactions between atoms in molecules  [^1] Invariance under translation is trivially satisfied by using relative positions as inputs..
Spherical Harmonics and Steerable Vectors
Spherical harmonics, a set of orthonormal basis functions defined over the sphere $\mathbb{S}^{2}$, are commonly employed in equivariant models. The real-valued spherical harmonics are typically denoted as $\{Y^{(l,m)}:\mathbb{S}^{2}\to\mathbb{R}\}$, where $l$ represents the degree and $m$ the order. For any orientation $\vec{\mathbf{p}}$, we define $\mathbf{Y}^{(l)}(\vec{\mathbf{p}})=[Y^{(l,-l)}(\vec{\mathbf{p}}),Y^{(l,-l+1)}(\vec{\mathbf{p}}),...,Y^{(l,l)}(\vec{\mathbf{p}})]$, a vector of size $2l+1$, and $\mathbf{Y}^{(0\to l)}(\vec{\mathbf{p}})=[\mathbf{Y}^{(0)}(\vec{\mathbf{p}}),...,\mathbf{Y}^{(l)}(\vec{\mathbf{p}})]$, a vector of size $(l+1)^{2}$.
A key property of spherical harmonics is their behavior under rotation $\mathbf{R}\in SO(3)$:
$\mathbf{Y}^{(l)}(\mathbf{R}\vec{\mathbf{p}})=\mathbf{D}^{(l)}(\mathbf{R})\mathbf{Y}^{(l)}(\vec{\mathbf{p}}),$ (3)
where $\mathbf{D}^{(l)}(\mathbf{R})$ is a $(2l+1)\times(2l+1)$ matrix known as the Wigner-D matrix of degree $l$. Notably, $\mathbf{D}^{(1)}(\mathbf{R})=\mathbf{R}$. Thus, $\mathbf{R}$ and $\mathbf{D}^{(l)}(\mathbf{R})$ can represent $T(g)$ and $T^{\prime}(g)$ in Equation (2). Following the convention in (chami2019hyperbolic, ; brandstetter2021geometric, ), we say that $\mathbf{Y}^{(l)}(\vec{\mathbf{p}})$ is steerable by the Wigner-D matrix of the same degree $l$. Furthermore, a vector that transforms according to an $l$-degree Wigner-D matrix is termed an $l$-degree steerable vector or a type-$l$ vector, residing in the vector space $\mathbb{V}_{l}$. Further mathematical details are available in Appendix A.1.
A common practice in equivariant models is to use steerable vectors as node embedding and encode relative positions $\vec{\mathbf{r}}_{ij}$ using spherical harmonics as edge attributes. TFN thomas2018tensor and NequIP batzner2022 provide a general framework that learns the interaction between node embeddings and edge attributes through equivariant convolution filters. These filters are composed of equivariant operations such as degree-wise linear (DW-Linear) layers, Clebsch-Gordan (CG) tensor products, and Gate mechanisms. SEGNN (brandstetter2021geometric, ) extends this convolution framework by more non-linear operations to achieve enhanced learning ability. SE(3)-Transformer fuchs2020se and Equiformer liao2023equiformer introduce an attention mechanism that assigns rotation-invariant weights to edge attributes, thereby improving the learning of key atomic interactions. Nevertheless, the core operation in these works remains the convolution filter based on the CG tensor product. Additionally, GotenNet aykentgotennet proposes an effective and lightweight structure that separates steerable vectors into an $l=0$ invariant component and $l>0$ equivariant components, subsequently interacting these parts using inner products and scalar multiplication. In summary, due to the requirement of equivariance, most equivariant model architectures are constrained to specific equivariant operations. In this work, we transform the steerable vector into spherical spatial domain, enabling the application of more flexible and highly non-linear structures within message passing.
Spherical Fourier Transform
Any square-integrable function $f(\vec{\mathbf{p}})$ defined over the sphere $\mathbb{S}^{2}$ can be expressed in a spherical harmonic basis via the spherical Fourier transform $\mathcal{F}$:
$f(\vec{\mathbf{p}})=\mathcal{F}(\mathbf{x})=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}x^{(l,m)}Y^{(l,m)}(\vec{\mathbf{p}}),$ (4)
where $x^{(l,m)}$ is the spherical Fourier coefficient. Conversely, the coefficient can be obtained from the function $f(\vec{\mathbf{p}})$ via the inverse transform $\mathcal{F}^{-1}$:
$x^{(l,m)}=\mathcal{F}^{-1}(f(\vec{\mathbf{p}}))=\int_{\Omega}f(\vec{\mathbf{p}})Y^{(l,m)*}(\vec{\mathbf{p}})d\Omega,$ (5)
where $d\Omega$ denotes the unit sphere and $Y^{(l,m)*}(\cdot)$ is the conjugate of the spherical harmonic function. The function $f(\vec{\mathbf{p}})$ and its coefficients $\mathbf{x}$ are referred to as the representations in the spherical spatial domain and the harmonic domain, respectively. SCN zitnick2022spherical introduced an approach that applies a simple point-wise neural network in the spherical spatial domain. This technique has been subsequently adopted by several works passaro2023reducing ; liao2024equiformerv within their message-passing frameworks. However, point-wise neural networks inherently limit their expressiveness. Our work aims to extend the capabilities of learning within the whole spherical spatial domain.
3 Model
In this section, we first make a review of tensor product to detail our motivation. Then, we introduction the formulation of our EST and clarify its properties.
3.1 Review of Clebsch-Gordan Tensor Product
A common approach in equivariant models is to incorporate steerable vectors with multiple degrees for node or edge features and then employing CG tensor products $\otimes:\mathbb{V}_{l_{1}}\times\mathbb{V}_{l_{2}}\to\mathbb{V}_{l}$, which is defined as:
$(\mathbf{x}_{1}\otimes\mathbf{x}_{2})^{(l,m)}=\sum^{l_{1}}_{m_{1}=-l_{1}}\sum^{l_{2}}_{m_{2}=-l_{2}}\mathcal{C}^{(l,m)}_{(l_{1},m_{1})(l_{2},m_{2})}\mathbf{x}^{(l_{1},m_{1})}_{1}\mathbf{x}^{(l_{2},m_{2})}_{2}$, (6)
where $\mathcal{C}^{(l,m)}_{(l_{1},m_{1})(l_{2},m_{2})}$ are the CG coefficients, a sparse tensor yielding non-zero terms when $|l_{1}-l_{2}|\leq l\leq(l_{1}+l_{2})$. These products enable interactions between various combinations of $l_{1}$ and $l_{2}$, crucial for the expressive power of MPNNs in representing latent equivariant functions dym2021on ; joshi2023expressive . However, CG tensor products lack high non-linearities brandstetter2021geometric , often requiring the stacking of multiple layers to capture complex molecular features. Furthermore, their high computational complexity ($\mathcal{O}(L^{6})$) luo2024enabling makes architectures with stacked tensor product layers computationally expensive brandstetter2021geometric . While techniques like Gate mechanisms brandstetter2021geometric ; liao2023equiformer and geometry-aware tensor attention aykent2025gotennet can be used to mitigate this limitation, they apply non-linear transformations to the invariant ($l=0$) representations. Higher-degree representations ($l>0$) are then updated through a scaling multiplication:
$C_{inv}=\mathrm{NL}(\mathbf{x}^{(0)},\sum^{L}_{l=1}\mathrm{d}(\mathbf{x}^{(l)})),\quad\text{and}\quad\tilde{\mathbf{x}}^{(l)}=C_{inv}\cdot\mathbf{x}^{(l)},$ (7)
where $C_{inv}$ is an invariant scalar, $\mathrm{NL}(\cdot)$ is a network with non-linear functions, and $\mathrm{d}(\cdot):\mathbb{V}_{l}\rightarrow\mathbb{V}_{0}$ maps equivariant features to invariants, such as an inner product. This indirect exchange of information between higher-degree representations still limits the overall expressiveness.
3.2 Equivariant Spherical Transformer
3.2.1 Node embedding
Steerable Representation
For representation with steerable vectors, we first define a maximal degree $L$. Each node $n$’s embedding $\mathbf{x}_{n}$ comprises $C$ channels, and each channel is a concatenation of steerable vectors from degree 0 to $L$: $\mathbf{x}_{n,c}=[x^{(0)}_{n,c},\mathbf{x}^{(1)}_{n,c},...,\mathbf{x}^{(L)}_{n,c}$]. As a result, the dimension of a steerable node embedding is $(L+1)^{2}\times C$.
Spatial Representation.
Steerable representations $\mathbf{x}_{n,c}$ can be transformed into a spherical function $f_{n,c}(\vec{\mathbf{p}})$ via Equation (4), with the summation truncated at a maximum degree $L$. Furthermore, $f_{n,c}(\vec{\mathbf{p}})$ is discretely represented by sampling $S$ points on the sphere, yielding a spatial node embedding $\mathbf{f}^{*}_{n}\in\mathbb{R}^{S\times C}$. It can be expressed as the concatenation over sampled points:
$\mathbf{f}^{*}_{n,c}=[(\mathbf{x}_{n,c}^{(0\to L)})^{T}\mathbf{Y}^{(0\to L)}(\vec{\mathbf{p}}_{1}),(\mathbf{x}_{n,c}^{(0\to L)})^{T}\mathbf{Y}^{(0\to L)}(\vec{\mathbf{p}}_{2}),...,(\mathbf{x}_{n,c}^{(0\to L)})^{T}\mathbf{Y}^{(0\to L)}(\vec{\mathbf{p}}_{S})].$ (8)
We use the term $\mathbf{f}^{*}_{n,c,s}$ to denote the signal at the sampled point $\vec{\mathbf{p}}_{s}$ of channel $c$ of node $n$. The inverse transform in Equation (5) can convert $\mathbf{f}^{*}_{n}$ back to $\mathbf{x}_{n}$.
In our model, the standard state of the node embedding is a steerable representation and it is transformed into a spatial representation when specific spherical operations are performed. Furthermore, the initial node embedding is constructed from two components: invariant mappings of the atomic attributes (the $l=0$ part) and the spherical harmonics of its neighbors’ relative positions (the $l>0$ part).
Operations in spatial domain
[…]
3.2.2 Spherical Attention
Treating $\mathbf{f}^{*}$ as a sequential data, where each sampled point $s$ has a $C$-dimensional feature vector, we apply a attention mechanism vaswani2017attention :
$a_{s_{i},s_{j}}=\frac{\exp(\mathbf{Q}_{s_{i}}\mathbf{K}_{s_{j}}^{T}/\sqrt{C})}{\sum_{s_{k}=1}^{S}\exp(\mathbf{Q}_{s_{i}}\mathbf{K}_{s_{k}}^{T}/\sqrt{C})},\quad\text{and}\quad\tilde{\mathbf{f}}^{*}_{s_{i}}=\sum_{s_{j}=1}^{S}a_{s_{i},s_{j}}\mathbf{V}_{s_{j}},$ (9)
where the $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ matrices are obtained through point-wise linear (PW-Linear) transformations of $\mathbf{f}^{*}$. The spherical attention (SA) is the core of EST. We also incorporate the pre-LayerNorm strategy xiong2020layer and feedforward networks (FFN) in EST. Our EST is applied in two ways: (i) for interactions between two spatial representations, $\mathbf{Q}$ is derived from one, and $\mathbf{K}$ and $\mathbf{V}$ from the other; (ii) for learning features within a single spatial representation, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are all derived from it.
Expressiveness
Assuming no information loss during the spherical Fourier transform and its inverse, i.e., $\mathbf{x}=\mathcal{F}^{-1}\mathcal{F}(\mathbf{x})$, Theorem 1 can evaluate the expressive power of EST.
Theorem 1.
For any two steerable representations $\mathbf{u}\in\mathbb{V}_{0\to l_{1}}$ and $\mathbf{v}\in\mathbb{V}_{0\to l_{2}}$, the spatial representation of their CG tensor product, $\mathbf{u}\otimes\mathbf{v}$, can be uniformly approximated by EST modules operating on the spatial representations of $\mathbf{u}$ and $\mathbf{v}$.
The proof can be found in Appendix B.1. We note that a Fourier transform without information loss is attainable by satisfying the Nyquist sampling rate, which mandates a minimum of $(2L)^{2}$ sampling points on the sphere zitnick2022spherical . Therefore, Theorem 1 substantiates the potential for substituting tensor products with EST. Furthermore, we present an empirical comparison of the expressiveness of EST and tensor product-based operations in Section 4.3, where our EST significantly surpasses the theoretical upper bound of tensor products.
Equivariance
We demonstrate that the equivariance of EST is related to the uniformity of the sampling points in Fourier transform.
Theorem 2.
If the sampling point density is the same across any local region of the sphere, EST operation preserves SO(3)-equivariance:
$\mathcal{F}^{-1}\big{(}\mathrm{EST}\big{(}\mathcal{F}(\mathbf{Dx})\big{)}\big{)}=\mathbf{D}\mathcal{F}^{-1}\big{(}\mathrm{EST}\big{(}\mathcal{F}(\mathbf{x})\big{)}\big{)},$ (10)
where $\mathbf{D}$ represents an arbitrary Wigner-D rotation matrix.
The proof is provided in Appendix B.2. Most prior works passaro2023reducing ; liao2024equiformerv implement spherical Fourier transforms using the e3nn library e3nn . However, e3nn does not guarantee a uniform sampling point density across all regions of the sphere. To address this, we redefine the sampling implementation using Fibonacci Lattices (FL) gonzalez2010measurement . The Cartesian coordinate of each sampling point $s$ is defined by
$\vec{\mathbf{p}}_{s}=[\sqrt{1-z_{s}^{2}}\cos(2s\pi/\lambda),\sqrt{1-z_{s}^{2}}\sin(2s\pi/\lambda),z_{s}],$ (11)
where $z_{s}=\frac{2s-1}{S-1}$ and $\lambda=\frac{1+\sqrt{5}}{2}$ denotes the golden ratio. The Fourier basis functions must satisfy orthogonality, i.e., $\int_{\Omega}Y^{(l,m)}(\vec{\mathbf{p}})Y^{(l^{\prime},m^{\prime})*}(\vec{\mathbf{p}})d\Omega=\delta_{ll^{\prime}}\delta_{mm^{\prime}}$, where $\delta_{ij}$ is the Kronecker delta. Consequently, we redefine the implementation of conjugate functions for the inverse transform as:
$Y^{(l,m)*}(\vec{\mathbf{p}})=\lambda^{(l,m)}Y^{(l,m)}(\vec{\mathbf{p}}),$ (12)
where $\lambda^{(l,m)}=1/\sum_{s}|Y^{(l,m)}(\vec{\mathbf{p}}_{s})|^{2}$. We provide detailed discussions between our sampling and the e3nn implementation in the Appendix B.2. In addition, we conducted ablation experiments in Section 4.3 to demonstrate that sampling methods with poor uniformity can severely compromise equivariance.
Relative Orientation Embedding on the Sphere
Despite its strong expressiveness, EST lacks the inherent ability to capture orientation relationships between sampling points, potentially leading to ambiguities dufter2022position . For example, if the features at points $s_{1}$ and $s_{2}$ are identical ($\mathbf{f}^{*}_{s_{1}}=\mathbf{f}^{*}_{s_{2}}$), their contributions to other sampling points $s_{i}$ become indistinguishable. To address this, we introduce a relative orientation embedding by augmenting the query $\mathbf{Q}$ and key $\mathbf{K}$ vectors with orientation information:
$\mathbf{Q}_{s_{i}}:=[\mathbf{Q}_{s_{i}},\vec{\mathbf{p}}_{s_{i}}],\quad\mathbf{K}_{s_{j}}:=[\mathbf{K}_{s_{j}},\vec{\mathbf{p}}_{s_{j}}],\quad\forall s_{i},s_{j}\in\{1,2,...,S\}.$ (13)
Consequently, the inner product between $\mathbf{Q}_{s_{i}}$ and $\mathbf{K}_{s_{j}}$ now incorporates a term aware of the orientation, related to $\vec{\mathbf{p}}^{T}_{s_{i}}\vec{\mathbf{p}}_{s_{j}}$. Moreover, this augmentation does not compromise the equivariant inductive bias of EST, as proven in Appendix B.3. The intuition behind our approach shares similarity with rotary position embeddings in NLP su2024roformer , where relative position embedding should be invariant under transformation. However, while NLP focuses on permutation invariance, our approach addresses rotation invariance, necessitating a fundamentally different implementation.
Figure 2: The architecture and building blocks of EST. SH and DTP denote spherical harmonic embedding and depth-wise tensor product liao2023equiformer , respectively. For simplicity, the Fourier and inverse Fourier transform steps are omitted. (a) Overall architecture. (b) Message block. (c) Update block. (d) Two experts operating on the steerable and spatial representations, respectively.
3.2.3 Mixture of Hybrid Experts
To enhance the model’s representational capacity, we utilize equivariant FFNs within a Mixture of Experts (MoE) framework shazeer2017outrageously . Starting with the steerable representation $\mathbf{x}$, the invariant ($l=0$) component is used to compute expert weights through a gate function:
$G(\mathbf{x})=\mathrm{Softmax}(\mathbf{x}^{(0)}\mathbf{W}_{G}),$ (14)
where $\mathbf{W}_{G}\in\mathbb{R}^{C\times E}$ is a trainable weight matrix. As illustrated in Figure 2(c), we employ two distinct expert structures: (1) a steerable FFN, operating on steerable inputs and consisting of two DW-Linear layers with an intermediate Gate activation liao2023equiformer ; and (2) a spherical FFN, processing spatial inputs and comprising two PW-Linear layers with an intermediate SiLU activation elfwing2018sigmoid . The final output is a weighted combination of the outputs from these hybrid experts, with the weights determined by Equation (14). These two expert types offer a crucial trade-off: the steerable FFN guarantees strict equivariance, while the spherical FFN provides enhanced expressive power. We therefore adapt their numbers based on the specific demands of the task. Additionally, we investigate common MoE techniques, including balancing loss lepikhin2020gshard and shared experts dai2024deepseekmoe . Further details and ablation studies can be found in Appendix C.
3.3 Overall Architecture
EST can be employed both in the message block to compute edge features and in the update block to refine node embeddings. Figure 2(b,c) defines a message passing layer with EST. When computing the message, the $\mathbf{Q}$ in SA come from the aggregation of node $i$ and $j$ embeddings, while $\mathbf{K}$ and $\mathbf{V}$ is derived from the spherical harmonic representation of the relative position $\vec{\mathbf{r}}_{ij}$. In the SA of the update block, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are all derived from the aggregated message.
EST serves as a flexible building block and can also be integrated into other equivariant models. Since the primary function of EST is to extract critical interactions or geometric features from complex high-degree representations, we prefer to replace the update blocks in existing models with our EST-based update block, where the input (aggregated message) contains sufficient and minimally processed neighbor features. With this strategy, EST-based architectures can also benefit from the computational efficiency of lightweight message blocks, while keep advantages to enhance the expressiveness.
4 Experiments
In this section, we construct experiments to investigate the efficacy of the proposed method. We evaluate its performance on the S2EF/IS2RE tasks from the OC20 chanussot2021open benchmark and QM9 ramakrishnan2014quantum tasks. Our analysis involves a comparison with a wide range of baseline models, including Schnet schutt2018schnet , PaiNN schutt2021equivariant , SEGNN brandstetter2021geometric , TFN thomas2018tensor , Dimenet++ klicpera2020fast , Equiformer liao2023equiformer , and Equiformerv2 liao2024equiformerv . Specifically for the OC20 benchmark, we extend our comparison to include SpinConv shuaibi2021rotation , GemNet gasteiger2021gemnet ; gasteiger2022gemnetoc , SphereNet liu2021spherical , SCN zitnick2022spherical and eSCN passaro2023reducing . Similarly, our QM9 experiments are augmented with comparisons to TorchMD-NET tholke2022equivariant , EQGAT le2022equivariant and GotenNet aykentgotennet . The specific configurations employed for each baseline model can be found in Appendix D.1.
Table 1: S2EF results of models trained on S2EF-All training dataset. $\lambda_{E}$ is the coefficient of the energy loss. “All+MD” denotes training models with additional OC20 MD dataset.
Number of Throughput Energy MAE Force MAE
Model parameters Samples/s $\mathrm{(}\mathrm{m}\mathrm{eV}\mathrm{)}$ $\downarrow$ $\mathrm{(}\mathrm{m}\mathrm{eV}\mathrm{/}\mathrm{\SIUnitSymbolAngstrom}\mathrm{)}$ $\downarrow$
SchNet 9.1M - $548.525$ $56.8$
DimeNet++-L-F+E 10.7M 4.6 $515.475$ $32.75$
SpinConv 8.5M 6.0 371.011 65371.011\,65 41.197 87541.197\,875
GemNet-dT 32M 25.8 315.118 775315.118\,775 27.172 47527.172\,475
GemNet-OC 39M 18.3 244.289 799 999 999 99244.289\,799\,999\,999\,99 21.719 050 000 000 00321.719\,050\,000\,000\,003
SCN (20 layers, test) 271M - $244$ $17.7$
eSCN (20 layers, test) 200M 2.9 $242$ $17.1$
EquiformerV2 ($\lambda_{E}=4$, 8 layers, All+MD) 31M 7.1 $232$ $16.26$
EquiformerV2 ($\lambda_{E}=2$, 20 layers) 153M 1.8 $236$ 15.7
EST ($\lambda_{E}=4$, 8 layers) 45M 6.8 231 $16.1$
4.1 OC20 Results
The OC20 dataset, comprising over 130 million molecular structures with force and energy labels, covers a broad spectrum of materials, surfaces, and adsorbates. We evaluate our models on two core sub-datasets of OC20: S2EF and IS2RE. All experimental configurations are detailed in Appendix D.2. Notably, We did not use very deep models or very long training schedules. For instance, our S2EF experiments utilize only 8 layers, and our IS2RE experiments employ 6 layers, significantly fewer than many advanced methods. Furthermore, as mentioned in Section 3.3, to enhance computational efficiency and facilitate a more direct comparison of different fundamental modules, we replace the EST architecture’s message module with the graph attention modules from EquiformerV2 and Equiformer for S2EF and IS2RE tasks, respectively.
S2EF Results
We trained an 8-layer EST model on the S2EF-All dataset and evaluated its performance on the S2EF validation sets. Each validation set was divided into four similarly-sized subsets: ID, OOD Ads, OOD Cat, and OOD Both. Consistent with prior work passaro2023reducing ; liao2024equiformerv , the best model was selected based on its performance on the ID validation subset during training and subsequently evaluated across all validation subsets. For comparison with SCN and eSCN, we report their test results due to the absence of publicly available validation results; previous work liao2024equiformerv indicates that the validation and test sets have a similar distribution. We compared our results with two variants of the SOTA EquiformerV2 model: a deep 20-layer architecture and one trained with additional MD dataset. As shown in Table 1, EST outperformed both EquiformerV2 variants in energy prediction. Furthermore, EST achieved competitive results in force prediction, surpassing several deeper baselines including SCN and eSCN.
Table 2: IS2RE results of models trained on IS2RE training dataset. Bold and underline indicate the best result, and the second best result, respectively.
Energy MAE (meV) $\downarrow$ EwT (%) $\uparrow$
Model ID OOD Ads OOD Cat OOD Both ID OOD Ads OOD Cat OOD Both
SchNet 639 734 662 704 2.96 2.33 2.94 2.21
PaiNN 575 783 604 743 3.46 1.97 3.46 2.28
TFN 584 766 636 700 4.32 2.51 4.55 2.66
DimeNet++ 562 725 576 661 4.25 2.07 4.10 2.41
GemNet-dT 527 758 549 702 4.59 2.09 4.47 2.28
GemNet-OC 560 711 576 671 4.15 2.29 3.85 2.28
SphereNet 563 703 571 638 4.47 2.29 4.09 2.41
SEGNN 533 692 537 679 5.37 2.46 4.91 2.63
Equiformer 504 688 521 630 5.14 2.41 4.67 2.69
SCN (16 layers) 516 643 530 604 4.92 2.71 4.42 2.76
EST (6 layers) 501 652 502 578 5.16 2.67 5.16 2.76
IS2RE Results
We trained our model directly on the IS2RE training set for energy prediction, without utilizing S2EF data. The results, summarized in Table 2, show that EST achieves SOTA performance on the ID, OOD Ads and OOD Both tasks and the second best on OOD Cat. Notably, the current SOTA models on OOD Cat, SCN, rely on complex 16-layer architectures with over 100M parameters. In contrast, EST employs a unified 6-layer structure comprising only 32.47M parameters, yet yields performance closely approaching theirs. Crucially, SCN relax equivariance to gain expressiveness, which can potentially lead to unstable predictions under input rotation. EST, conversely, provides more stable predictions owing to its stronger equivariance. Furthermore, in these IS2RE experiments, the EST architecture incorporates the message block from Equiformer. The key architectural difference resides in the update block. We observe that EST consistently surpasses the Equiformer model across all evaluated metrics, providing further evidence for the effectiveness of the proposed EST module and the model merging strategy discussed in Section 3.3.
Table 3: Results on QM9 dataset for various properties. $\dagger$} denotes using different data partitions.
Task $\alpha$ $\Delta\varepsilon$ $\varepsilon_{HOMO}$ $\varepsilon_{LUMO}$ $\mu$ $C_{v}$ $G$ $H$ $R^{2}$ $U$ $U_{0}$ ZPVEZPVE
Units $bohr^{3}$ meV meV meV D cal//(mol K) meV meV $bohr^{3}$ meV meV meV
SchNet .235 63 41 34 .033 .033 14 14 .073 19 14 1.70
TFN .223 58 40 38 .064 .101 - - - - - -
DimeNet++ .044 33 25 20 .030 .023 8 7 .331 6 6 1.21
PaiNN .045 46 28 20 .012 .024 7.35 5.98 .066 5.83 5.85 1.28
TorchMD-NET .059 36 20 18 .011 .026 7.62 6.16 .033 6.38 6.15 1.84
SEGNN .060 42 24 21 .023 .031 15 16 .660 13 15 1.62
EQGAT .053 32 20 16 .011 .024 23 24 .382 25 25 2.00
Equiformer .046 30 15 14 .011 .023 7.63 6.63 .251 6.74 6.59 1.26
EquiformerV2 .050 29 14 13 .010 .023 7.57 6.22 .186 6.49 6.17 1.47
EST .042 28 13 12 .011 .022 7.03 5.94 .298 5.92 5.64 1.31
EST (with GA) .041 29 14 13 .011 .021 7.18 6.17 .227 6.35 6.32 1.27
4.2 QM9 Results
The QM9 benchmark comprises quantum chemical properties for a relevant, consistent, and comprehensive chemical space of 134k equilibrium small organic molecules containing up to 29 atoms. Each atom is represented by its 3D coordinates and an embedding of its atomic type (H, C, N, O, F). We developed two model architectures: a 6-layer model employing the message-passing layer detailed in Figure 2, and a separate model that integrates message blocks (GA module) from Equiformer. For additional configuration specifics, please refer to Appendix D.2.
Given that the QM9 dataset is considerably smaller than OC20, models are more prone to overfitting if equivariance is destroied. Ablation study in Table 5 shows that practical implementation for Fourier transform may lead to minor deviations from perfect equivariance. Nevertheless, EST and EST (with GA) achieve best on eight of the twelve tasks (see Table 3). Notably, EST (with GA) significantly outperformed Equiformer despite utilizing a similar message block. Furthermore, we included a comparison with GotenNet aykent2025gotennet , which employs a more extensive training schedule. To ensure a fair assessment, we applied this identical schedule to EST as well; see Appendix D.3 for details.
4.3 Ablation study
In this section, we explore two key properties of EST: expressiveness and equivariance. The other experiments can be found in Appendix D, where we comprehensively investigate the influence of building components, including SA, spherical relative orientation embedding and hybrid experts.
Table 4: Experiments on Rotationally symmetric structures.
GNN Layer 2 fold 3 fold 10 fold 100 fold GNN Layer 2 fold 3 fold 5 fold 10 fold
ESTL=1 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 TFN/MACEL=1 50.0 ± 0.0 50.0 ± 0.0 50.0 ± 0.0 50.0 ± 0.0
ESTL=2 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 TFN/MACEL=2 100.0 ± 0.0 50.0 ± 0.0 50.0 ± 0.0 50.0 ± 0.0
ESTL=3 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 TFN/MACEL=3 100.0 ± 0.0 100.0 ± 0.0 50.0 ± 0.0 50.0 ± 0.0
ESTL=5 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 TFN/MACEL=5 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 50.0 ± 0.0
ESTL=10 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 TFN/MACEL=10 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0 100.0 ± 0.0
Expressiveness
Following joshi2023expressive , we employ evaluation metrics that distinguish $n$-fold symmetric structures to precisely assess EST’s expressive power. We first construct a single-layer message-passing layer based on Figure 2, consistent with tensor product-based operations (TFN, MACE). Additionally, we remove the steerable FFN for a clear comparison. As demonstrated in dym2021on ; joshi2023expressive , the expressive power of tensor product-based models is limited by the maximum degree $L$, failing to perfectly distinguish $n$-fold symmetric structures when $n>L$. From Table 4, we observe that EST significantly breaks through this theoretical boundary. It allows the model to distinguish symmetric structures even with very high fold (e.g., using 1-degree EST for 100-fold structures).
Table 5: Equivariance evaluation with different sampling strategies.
Sampling Strategy/Number 1-layer 2-layer 3-layer 4-layer 5-layer 6-layer
FL / $[64,128]$ 0.0009 0.0013 0.0014 0.0016 0.0025 0.0018
FL / $[128,256]$ 0.0004 0.0010 0.0007 0.0005 0.0015 0.0013
FL / $[256,256]$ 0.0002 0.0002 0.0002 0.0003 0.0001 0.0002
e3nn / $[210,210]$ 0.0084 0.1006 0.0366 0.0593 0.0046 0.0199
Equivariance
While we theoretically demonstrate that EST can achieve strict equivariance, perfect uniform sampling is challenging in practice, potentially leading to loss of equivariance. To investigate it, we build untrained networks with 1 to 6 layers based on Figure 2 and control the number of spherical sampling points in the Fourier transforms within the message and update blocks. We randomly select 1000 molecules from QM9, compute their outputs $y_{1},...,y_{1000}$, and then compute the outputs after applying random rotations, $\hat{y}_{1},...,\hat{y}_{1000}$. The average absolute error $(1/1000)\sum_{i=1}^{1000}|y_{i}-\hat{y}_{i}|$ serves as our measure of equivariance loss. As shown in Table 5, EST with FL sampling closely approximates strict equivariance even without training, and its equivariance further improves with an increasing number of spherical sampling points. Additionally, EST with the sampling implementation in e3nn show a poor equivariance.
5 Conclusion
We present EST, a new SE(3)-equivariant framework for modeling molecules. By integrating a Transformer structure with steerable vectors, EST offers greater expressive power than tensor-product-based frameworks. We showed EST’s strong performance on the OC20 and QM9 datasets. A current limitation is the equivariance loss caused by spherical sampling. Future work will focus on improving spherical sampling uniformity. This could lead to fewer sampling points and even better equivariance.
APPENDIX
[…]
$\printcontents[Appendix]$l1
Appendix A The Mathematics
A.1 The Mathematics of Spherical Harmonics
A.1.1 The Properties of Spherical Harmonics
The spherical harmonics $Y^{(l,m)}(\theta,\varphi)$ are the angular portion of the solution to Laplace’s equation in spherical coordinates where azimuthal symmetry is not present. Some care must be taken in identifying the notational convention being used. In this entry, $\theta$ is taken as the polar (colatitudinal) coordinate with $\theta$ in $0,\pi]$, and $\varphi$ as the azimuthal (longitudinal) coordinate with $\varphi$ in $0,2\pi)$.
Spherical harmonics satisfy the spherical harmonic differential equation, which is given by the angular part of Laplace’s equation in spherical coordinates. If we define the solution of Laplace’s equation as $F=\Phi(\varphi)\Theta(\theta)$, the equation can be transformed as:
$\frac{\Phi(\varphi)}{\sin\theta}\frac{d}{d\theta}\left(\sin\theta\frac{d\Theta}{d\theta}\right)+\frac{\Theta(\theta)}{\sin^{2}\theta}\frac{d^{2}\Phi(\varphi)}{d\varphi^{2}}+l(l+1)\Theta(\theta)\Phi(\varphi)=0$ (15)
Here we omit the derivation process and just show the result. The (complex-value) spherical harmonics are defined by:
$Y^{(l,m)}(\theta,\varphi)\equiv\sqrt{\frac{2l+1}{4\pi}\frac{(l-m)!}{(l+m)!}}P^{(l,m)}(\cos\theta)e^{im\varphi},$ (16)
where $P^{(l,m)}(\cos\theta)$ is an associated Legendre polynomial. Spherical harmonics are integral basis, which satisfy:
$\begin{array}[]{l}\int_{0}^{2\pi}\int_{0}^{\pi}Y^{(l_{1},m_{1})}(\theta,\varphi)Y^{(l_{2},m_{2})}(\theta,\varphi)Y^{(l_{3},m_{3})}(\theta,\varphi)\sin\theta d\theta d\varphi\\ =\sqrt{\frac{\left(2l_{1}+1\right)\left(2l_{2}+1\right)\left(2l_{3}+1\right)}{4\pi}}\left(\begin{array}[]{ccc}l_{1}&l_{2}&l_{3}\\ 0&0&0\end{array}\right)\left(\begin{array}[]{ccc}l_{1}&l_{2}&l_{3}\\ m_{1}&m_{2}&m_{3}\end{array}\right),\end{array}$ (17)
where $\left(\begin{array}[]{ccc}l_{1}&l_{2}&l_{3}\\ m_{1}&m_{2}&m_{3}\end{array}\right)$ is a Wigner 3j-symbol (which is related to the Clebsch-Gordan coefficients). We list a few spherical harmonics which are:
$Y^{(0,0)}(\theta,\varphi)$ $=\frac{1}{2}\sqrt{\frac{1}{\pi}},$ (18)
$Y^{(1,-1)}(\theta,\varphi)$ $=\frac{1}{2}\sqrt{\frac{3}{2\pi}}\sin\theta e^{-i\varphi},$
$Y^{(1,0)}(\theta,\varphi)$ $=\frac{1}{2}\sqrt{\frac{3}{\pi}}\cos\theta,$
$Y^{(1,1)}(\theta,\varphi)$ $=\frac{-1}{2}\sqrt{\frac{3}{2\pi}}\sin\theta e^{i\varphi},$
$Y^{(2,-2)}(\theta,\varphi)$ $=\frac{1}{4}\sqrt{\frac{15}{2\pi}}\sin{}^{2}\theta e^{-2i\varphi},$
$Y^{(2,-1)}(\theta,\varphi)$ $=\frac{1}{2}\sqrt{\frac{15}{2\pi}}\sin\theta\cos\theta e^{-i\varphi},$
$Y^{(2,0)}(\theta,\varphi)$ $=\frac{1}{4}\sqrt{\frac{5}{\pi}}\left(3\cos^{2}\theta-1\right),$
$Y^{(2,1)}(\theta,\varphi)$ $=\frac{-1}{2}\sqrt{\frac{15}{2\pi}}\sin\theta\cos\theta e^{i\varphi},$
$Y^{(2,2)}(\theta,\varphi)$ $=\frac{1}{4}\sqrt{\frac{15}{2\pi}}\sin{}^{2}\theta e^{2i\varphi},$
In this work, we use the real-value spherical harmonics rather than the complex-value one, which can be written as :
$Y^{0,0}(\theta,\varphi)$ $=\sqrt{\frac{1}{4\pi}},$ (19)
$Y^{(1,-1)}(\theta,\varphi)$ $=\sqrt{\frac{3}{4\pi}}\sin\varphi\sin\theta,$
$Y^{(1,0)}(\theta,\varphi)$ $=\sqrt{\frac{3}{4\pi}}\cos\theta,$
$Y^{(1,1)}(\theta,\varphi)$ $=\sqrt{\frac{3}{4\pi}}\cos\varphi\sin\theta,$
$Y^{(2,-2)}(\theta,\varphi)$ $=\sqrt{\frac{15}{16\pi}}\sin(2\varphi)\sin^{2}\theta,$
$Y^{(2,-1)}(\theta,\varphi)$ $=\sqrt{\frac{15}{4\pi}}\sin\varphi\sin\theta\cos\theta,$
$Y^{(2,0)}(\theta,\varphi)$ $=\sqrt{\frac{5}{16\pi}}(3\cos^{2}\theta-1),$
$Y^{(2,1)}(\theta,\varphi)$ $=\sqrt{\frac{15}{4\pi}}\cos\varphi\sin\theta\cos\theta,$
$Y^{(2,2)}(\theta,\varphi)$ $=\sqrt{\frac{15}{16\pi}}\cos(2\varphi)\sin^{2}\theta.$
A.1.2 Fourier transformation over $\mathbb{S}^{2}$
In the main paper, we show that any square-integrable function $f(\cdot)$ can thus be expanded as a linear combination of spherical harmonics:
$f(\vec{\mathbf{p}})=\sum_{l=0}^{\infty}\sum^{l}_{m=-l}\mathbf{x}^{(l,m)}Y^{(l,m)}(\vec{\mathbf{p}}),$ (20)
where $\vec{\mathbf{p}}=(\theta,\varphi)$ denotes the orientations, like what we do in the main paper. The coefficient $\mathbf{x}^{(l,m)}$ can be obtained by the inverse transformation over $\mathbb{S}^{2}$, which is
$\mathbf{x}^{l,m}=\int_{\Omega}f(\vec{\mathbf{p}})Y^{(l,m)*}(\vec{\mathbf{p}})d\Omega=\int_{0}^{2\pi}d\varphi\int_{0}^{\pi}d\theta\sin\theta f(\vec{\mathbf{p}})Y^{(l,m)*}(\vec{\mathbf{p}}).$ (21)
Using the fact $\mathbf{Y}^{l}(\mathbf{R}\vec{\mathbf{p}})=\mathbf{D}^{l}(\mathbf{R})\mathbf{Y}^{l}(\vec{\mathbf{p}}),$ and Equation (20), we know
$f(\mathbf{R}\vec{\mathbf{p}})=\sum_{l=0}^{\infty}\sum^{l}_{m=-l}\mathbf{x}^{(l,m)}Y^{(l,m)}(\mathbf{R}\vec{\mathbf{p}})=\sum_{l=0}^{\infty}\sum^{l}_{m=-l}\mathbf{x}^{(l,m)}\mathbf{D}Y^{(l,m)}(\vec{\mathbf{p}}).$ (22)
Therefore, we can get the conclusion that spatial representation $f(\mathbf{R}\vec{\mathbf{P}})$ and $f(\vec{\mathbf{P}})$ is steerable, which can be represented by
$f(\mathbf{R}\vec{\mathbf{p}})=\mathbf{D}^{-1}\mathbf{x}=\mathbf{D}^{T}\mathbf{x}.$ (23)
A.1.3 The Relationship Between Spherical Harmonics and Wigner-D Matrix
A rotation $\mathbf{R}$ sending the $\vec{\mathbf{p}}$ to $\mathbf{R}\vec{\mathbf{p}}$ can be regarded as a linear combination of spherical harmonics that are set to the same degree. The coefficients of linear combination represent the complex conjugate of an element of the Wigner D-matrix. The rotational behavior of the spherical harmonics is perhaps their quintessential feature from the viewpoint of group theory. The spherical harmonics $Y^{l,m}$ provide a basis set of functions for the irreducible representation of the group SO(3) with dimension $(2l+1)$.
The Wigner-D matrix can be constructed by spherical harmonics. Consider a transformation $Y^{l,m}(\vec{\mathbf{p}})=Y^{l,m}(\mathbf{R}_{\alpha,\beta,\gamma}\vec{\mathbf{p}}_{x})$, where $\vec{\mathbf{p}}_{x}$ denote the x-orientation. $\alpha,\beta,\gamma$ denotes the items of Euler angle. Therefore, $Y^{l,m}(\vec{\mathbf{p}})$ is invariant with respect to rotation angle $\gamma$. Based on this viewpoint, the Wigner-D matrix with shape $(2l+1)\times(2l+1)$ can be defined by:
$\mathbf{D}^{(l,m)}(\mathbf{R}_{\alpha,\beta,\gamma})=\sqrt{2l+1}Y^{(l,m)}(\vec{\mathbf{p}}).\\$ (24)
In this case, the orientations are encoded in spherical harmonics and their Wigner-D matrices, which are utilized in our cross module.
A.2 Equivariant Operation
A.2.1 Equivariance of Clebsch-Gordan Tensor Product
The Clebsch-Gordan Tensor Product shows a strict equivariance for different group representations, which make the mixture representations transformed equivariant based on Wigner-D matrices. We use $\mathbf{D}^{l,m^{\prime}m}$ to denote the element of Wigner-D matrix in the degree $l$. The Clebsch-Gordan coefficient satisfies:
[…]
Therefore, the spherical harmonics can be combined equivariantly by CG Tensor Product:
$CG\left(\sum_{m_{1}^{\prime}}\mathbf{D}^{(l_{1},m_{1}m_{1}^{\prime})}(g)Y^{(l_{1},m_{1}^{\prime})},\sum_{m_{2}^{\prime}}\mathbf{D}^{(l_{2},m_{2}m_{2}^{\prime})}(g)Y^{(l_{2},m_{2}^{\prime})}\right)$ (26)
$=$ $\sum_{m_{1},m_{2}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}\sum_{m_{1}^{\prime}}\mathbf{D}^{(l_{1},m_{1}m_{1}^{\prime})}(g)Y^{(l_{1},m_{1}^{\prime})}\sum_{m_{2}^{\prime}}\mathbf{D}^{(l_{2},m_{2}m_{2}^{\prime})}(g)Y^{(l_{2},m_{2}^{\prime})}$
$=$ $\sum_{m_{0}^{\prime}}\mathbf{D}^{(l_{0},m_{0}m_{0}^{\prime})}(g)\sum_{m_{1},m_{2}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0}m_{0}^{\prime}\right)}Y^{(l_{1},m_{1}^{\prime})}Y^{(l_{2},m_{2}^{\prime})}$
$=$ $\sum_{m_{0}^{\prime}}\mathbf{D}^{(l_{0},m_{0}m_{0}^{\prime})}(g)CG\left(Y^{(l_{1},m_{1}^{\prime})},Y^{(l_{2},m_{2}^{\prime})}\right).$
Here, we omit the input argument of the spherical harmonics, which can represent any direction on the sphere. Equation (26) represents a relationship between scalar. If we transform the scalar to vector or matrix like what we do in Equation (3), Equation (26) is equal to
$(\mathbf{D}^{l_{1}}\mathbf{u}\otimes\mathbf{D}^{l_{2}}\mathbf{v})^{l_{0}}=\mathbf{D}^{l_{0}}(\mathbf{u}\otimes\mathbf{v})^{l_{0}}.$ (27)
The tensor CG product mixes two representations to a new representation under special rule $|l_{1}-l_{2}|\leq l\leq(l_{1}+l_{2})$. For example, 1.two type-0 vectors will only generate a type-0 representations; 2.type-$l_{1}$ and type-$l_{2}$ can generate type-$l_{1}+l_{2}$ vector at most. Note that some widely-used products are related to tensor product: scalar product ($l_{1}=0$, $l_{2}=1$, $l=1$), dot product ($l_{1}=1$, $l_{2}=1$, $l=0$) and cross product ($l_{1}=1$, $l_{2}=1$, $l=1$). However, for each element with $l>0$, there are multi mathematical operation for the connection with weights. The relation between number of operations and degree is quadratic. Thus, as degree increases, the amount of computation increases significantly, making calculation of the CG tensor product slow for higher order irreps. This statement can be proven by the implementation of e3nn (o3.FullyConnectedTensorProduct).
A.2.2 Learnable Parameters in Tensor Product
Previous works utilize the e3nn library [28] to implement the corresponding tensor product. It is crucial to emphasize that the formulation of CG tensor product is devoid of any learnable parameters, as CG coefficients remain constant. In the context of e3nn, learnable parameters are introduced into each path, represented as $w(\mathbf{u}^{l_{1}}\otimes\mathbf{v}^{l_{2}})$. Importantly, these learnable parameters will not destory the equivariance of each path. However, they are limited in capturing directional information. In equivariant models, the original CG tensor product primarily captures directional information. We have previously mentioned our replacement of the CG tensor product with learnable modules. It is worth noting that our focus lies on the CG coefficients rather than the learnable parameters in the e3nn implementation.
A.2.3 Gate Activation and Normalization
Gate Activation. In equivariant models, the Gate activation combines two sets of group representations. The first set consists of scalar steerable vector ($l=0$), which are passed through standard activation functions such as sigmoid, ReLU and SiLU. The second set comprises higher-order steerable vector (($l>0$)), which are multiplied by an additional set of scalar steerable vector that are introduced solely for the purpose of the activation layer. These scalar steerable vector are also passed through activation functions.
Normalization. Normalization is a technique commonly used in neural networks to normalize the activations within each layer. It helps stabilize and accelerate the training process by reducing the internal covariate shift, which refers to the change in the distribution of layer inputs during training.
The normalization process involves computing the mean and variance across the channels. In equivariant normalization, the variance is computed using the root mean square value of the L2-norm of each type-$l$ vector. Additionally, this normalization removes the mean term. The normalized activations are then passed through a learnable affine transformation without a learnable bias, which enables the network to adjust the mean and variance based on the specific task requirements.
A.3 Relationship Between Expressive Power and Equivariant Operations
In [13], Theorem 2 establishes the universality of equivariant networks based on the TFN structure:
Theorem. For all $n\in\mathbb{N}$, $\mathbf{l}_{T}\in\mathbb{N_{+}}^{*}$,
  • 1. For $D\in\mathbb{N_{+}}$, every G-equivariant polynomial $p$ : $\mathbb{R}^{3\times n}$ $\to W_{\mathbf{l}_{T}}^{n}$ of degree $D$ is in $F^{TFN}_{C(D),D}$.
  • 2. Every continuous G-equivariant function can be approximated uniformly on compact sets by functions in $\cup_{D\in\mathbb{N_{+}}}F^{TFN}_{C(D),D}$.
Here, $n$ represents the number of input points (or nodes), $\mathbf{l}_{T}$ represents the degree of the approximated G-equivariant function, $C$ represents the number of channels, and $D$ represents the degree of the TFN (Tensor Field Network) structure, which is equivalent to the term $l$ used in our method. The TFN structure consists of two layers, including convolution and self-interaction. Self-interaction involves equivariant linear functions. The convolution operation calculates the CG tensor product between different steerable representations, which is a fundamental operation for transforming directional information. Most equivariant models based on group representations use a similar approach (CG tensor product) to capture directional features. Therefore, the theorem mentioned above also applies to building blocks based on CG tensor products, such as SEGNN [11] and Equiformer [12].
It is important to note that achieving an infinite degree in practice is not feasible. However, equivariant models based on group representations can enhance their expressive power by increasing the number of maximal degrees [13]. In their evaluation of expressive power, as presented in [14], the authors utilize the GWL (geometric Weisfeiler-Leman) graph isomorphism test. In Table 2 of their work, it is evident that equivariant models with a maximal degree denoted as $L$ are incapable of distinguishing $n$-fold symmetric structures when $n$ exceeds the value of $L$.
Appendix B Proofs and Details For Section 3
B.1 Proof for Expressiveness of EST
The proof for Theorem 1 is shown in the following.
Proof.
Given two steerable vectors $\mathbf{u}\in\mathbb{V}_{0\to l_{1}}$ and $\mathbf{v}\in\mathbb{V}_{0\to l_{2}}$, and we define the CG tensor product result is $\mathbf{w}\in\mathbb{V}_{l_{0}}$, where $l_{0}<=l_{1}+l_{2}$. The spatial representations of $\mathbf{w}$ can be represented as:
$\sum_{l_{0},m_{0}}\mathbf{w}^{(l_{0},m_{0})}Y^{(l_{0},m_{0})}=\sum_{l_{0},m_{0}}({\mathbf{u}^{(0\to l_{1})}}^{T}\mathbf{C}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}\mathbf{v}^{(0\to l_{2})})Y^{(l_{0},m_{0})},$ (28)
where $\mathbf{C}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$ is a matrix including the whole CG coefficients corresponding to degree $l_{0}$ and order $m_{0}$. We temporarily ignore the input of the spherical harmonics. The spherical Transformer use the multiplication between spatial representations of $\mathbf{u}$ and $\mathbf{v}$:
$\sum_{l_{1},m_{1}}\mathbf{u}^{(l_{1},m_{1})}Y^{(l_{1},m_{1})}\sum_{l_{2},m_{2}}\mathbf{v}^{(l_{2},m_{2})}Y^{(l_{2},m_{2})}$ (29)
$=$ $\sum_{l_{1},m_{1},l_{2},m_{2}}\mathbf{u}^{(l_{1},m_{1})}Y^{(l_{1},m_{1})}\mathbf{v}^{(l_{2},m_{2})}Y^{(l_{2},m_{2})}$
$=$ ${\mathbf{u}^{(0\to l_{1})}}^{T}({\mathbf{Y}^{(0\to l_{1})}}^{T}\mathbf{Y}^{(0\to l_{2})})\mathbf{v}^{(0\to l_{2})}$
Recall that CG coefficients are in fact the expansion coefficients of a product of two spherical harmonics in terms of a single spherical harmonic (see Equation (17)):
$Y^{(l_{1},m_{1})}Y^{(l_{2},m_{2})}=$ $\sum_{l_{0},m_{0}}\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)}{4\pi(2l_{0}+1)}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(0,0\right)}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}Y^{l_{0},m_{0}}$ (30)
Equation (LABEL:equ:mul_u_v) can be transformed to
${\mathbf{u}^{(0\to l_{1})}}^{T}({\mathbf{Y}^{(0\to l_{1})}}^{T}\mathbf{Y}^{(0\to l_{2})})\mathbf{v}^{(0\to l_{2})}$ (31)
$=$ $\sum_{l_{0},m_{0}}{\mathbf{u}^{(0\to l_{1})}}^{T}\mathbf{H}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}\mathbf{v}^{(0\to l_{2})}Y^{(l_{0},m_{0})},$
where $\mathbf{H}^{(l_{0},m_{0})}_{(l_{1},m_{1}),(l_{2},m_{2})}=\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)}{4\pi(2l_{0}+1)}}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(0,0\right)}\mathcal{C}_{\left(l_{1},m_{1}\right)\left(l_{2},m_{2}\right)}^{\left(l_{0},m_{0}\right)}$. Compared to $\mathbf{C}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$, the term $\mathbf{H}^{(l_{0},m_{0})}_{(0\to l_{1}),(0\to l_{2})}$ introduces an additional constant term that can be approximated by linear layers and the FFNs. Therefore, Equation (LABEL:equ:mul_u_v) is equivalent to Equation (28). On the other hand, Equation (LABEL:equ:mul_u_v) is inherently part of the Transformer architecture: when relationships between different orientations are masked and the $\mathbf{V}$ vectors are taken as constant vectors, the Transformer can naturally reduce to Equation (LABEL:equ:mul_u_v). ∎
Through the above proof, we find that the spherical representation of the output from the CG tensor product is fundamentally a special case of EST. Furthermore, EST can capture dependencies between different orientations, which is particularly beneficial for approximating higher degree spherical harmonic representations. For instance, at a direction $(\theta,\varphi)$, higher-degree spherical harmonics may involve contributions from other directions such as $(2\theta,\varphi),(\theta,2\varphi),(2\theta,2\varphi)$ (see $Y^{(2)}$ in Equation (19)). This property also enables EST to approximate equivariant features of higher degree than the input.
B.2 Proof for Equivariance of EST
The proof for Theorem 2 is shown in the following.
Proof.
Given a spatial representation $f(\vec{\mathbf{p}})$ transformed from a steerable representation $\mathbf{x}$, spherical attention can be represented as:
$\tilde{f}(\vec{\mathbf{p}}_{1})=\int_{\vec{\mathbf{p}}_{2}\in\Omega}a\big{(}f^{Q}(\vec{\mathbf{p}}_{1}),f^{K}(\vec{\mathbf{p}}_{2})\big{)}f^{V}(\vec{\mathbf{p}}_{2})d\vec{\mathbf{p}}_{2},$ (32)
where $a(\cdot)$ denotes the attention coefficients. Terms $Q,K,V$ denote three independent linear transformations. When the origin representation $\mathbf{x}$ is rotation by Wigner-D matrix $\mathbf{D}^{-1}$, the representation $f(\vec{\mathbf{p}})$ is transformed to $f(\mathbf{R}\vec{\mathbf{p}})$, where $\mathbf{R}$ and $\mathbf{D}$ share the same transformation parameters (see Equation (23)). Therefore, the attention results are changed to
$\int_{\mathbf{R}\vec{\mathbf{p}}_{2}\in\Omega}a\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1}),f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})\big{)}f^{V}(\mathbf{R}\vec{\mathbf{p}}_{2})d\mathbf{R}\vec{\mathbf{p}}_{2}$ (33)
$=$ $\int_{\mathbf{R}\vec{\mathbf{p}}_{2}\in\Omega}\frac{\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})\big{)}}{\int_{\mathbf{R}\vec{\mathbf{p}}_{3}\in\Omega}\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\mathbf{R}\vec{\mathbf{p}}_{3})\big{)}d\mathbf{R}\vec{\mathbf{p}}_{3}}f^{V}(\mathbf{R}\vec{\mathbf{p}}_{2})d\mathbf{R}\vec{\mathbf{p}}_{2}$
$=$ $\int_{\mathbf{R}\vec{\mathbf{p}}_{2}\in\Omega}\frac{\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})\big{)}}{\int_{\vec{\mathbf{p}}_{3}\in\Omega}\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\vec{\mathbf{p}}_{3})\big{)}d\vec{\mathbf{p}}_{3}}f^{V}(\mathbf{R}\vec{\mathbf{p}}_{2})d\mathbf{R}\vec{\mathbf{p}}_{2}$
$=$ $\int_{\vec{\mathbf{p}}_{2}\in\Omega}\frac{\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\vec{\mathbf{p}}_{2})\big{)}}{\int_{\vec{\mathbf{p}}_{3}\in\Omega}\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\vec{\mathbf{p}}_{3})\big{)}d\vec{\mathbf{p}}_{3}}f^{V}(\vec{\mathbf{p}}_{2})d\vec{\mathbf{p}}_{2}$
$=$ $\int_{\vec{\mathbf{p}}_{2}\in\Omega}a\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1}),f^{K}(\vec{\mathbf{p}}_{2})\big{)}f^{V}(\vec{\mathbf{p}}_{2})d\vec{\mathbf{p}}_{2}$
$=$ $\tilde{f}(\mathbf{R}\vec{\mathbf{p}}_{1}).$
Therefore, the output of spherical attention remains steerable after rotation. If we transform $\tilde{f}(\mathbf{R}\vec{\mathbf{p}}_{1})$ into its steerable representation, denoted as $\tilde{\mathbf{x}}$, the following relationship holds:
$\tilde{\mathbf{x}}=\mathbf{D}^{-1}\mathcal{F}^{-1}\big{(}\mathrm{SA}(\mathcal{F}(\mathbf{x}))\big{)}=\mathcal{F}^{-1}\big{(}\mathrm{SA}(\mathcal{F}(\mathbf{D}^{-1}\mathbf{x}))\big{)}$ (34)
The equivariance of spherical attention holds. Moreover, the spherical FFNs is obviously equivariant because the function $\tilde{f}(\vec{\mathbf{p}}_{1})=\mathrm{FFN}\big{(}f(\vec{\mathbf{p}}_{1})\big{)}$ only focus on one orientation. Therefore, we can prove that the whole EST framework contained spherical attention and spherical FFN is equivariant. ∎
In the implementation, the continuous function must be represented by discrete sampling. The attention operation become
$\tilde{f}(\vec{\mathbf{p}}_{1})=\sum_{\vec{\mathbf{p}}_{2}}a\big{(}f^{Q}(\vec{\mathbf{p}}_{1}),f^{K}(\vec{\mathbf{p}}_{2})\big{)}f^{V}(\vec{\mathbf{p}}_{2}).$ (35)
After introducing random rotations, Equation (35) becomes
$\tilde{f}(\mathbf{R}\vec{\mathbf{p}}_{1})=\sum_{\vec{\mathbf{p}}_{3}}a\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1}),f^{K}(\vec{\mathbf{p}}_{3})\big{)}f^{V}(\vec{\mathbf{p}}_{3}).$ (36)
Let us consider an ideal case where the point density is the same across all local regions. That is, the points on the sphere are perfectly uniformly distributed. In this scenario, the points in any region are symmetric, and we can find a set of $\vec{\mathbf{p}}_{3}$ that has a one-to-one correspondence with all $\mathbf{R}\vec{\mathbf{p}}_{2}$, i.e., the difference between the set of points $\vec{\mathbf{p}}_{2}$ and the set of points $\vec{\mathbf{p}}_{3}$ is only in their order. Therefore, spherical attention can still strictly preserve equivariance. In contrast, if we cannot find points that correspond one-to-one to all $\mathbf{R}\vec{\mathbf{p}}_{2}$, equivariance will be compromised. To reduce this equivariance loss, we can employ approximately uniform sampling methods, such as Fibonacci Lattices.
Sampling strategies
Most previous works use the e3nn implementation for spherical Fourier transform. However, it significantly destroy the uniformity of spherical sampling, which is illustrated in Figure 3(a). In contrast, Fibonacci Lattices (FL) do not directly divide the polar angle and azimuth angle into a grid. Instead, they select sampling points on the sphere in a spiral pattern. As shown in Figure 3(b), FL tends to achieve more uniform sampling, thereby improving the equivariance of EST. This is also consistent with the results observed in Table 5.
(a) e3nn Sampling
(b) Fibonacci Lattices Sampling
Figure 3: Two spherical sampling strategies.
B.3 Equivariance of Relative Orientation Embedding
The equivariance of relative orientation embedding can be proven with the same way in Equation (LABEL:equ:proof_equivariance), where $a\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1}),f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})\big{)}$ is transformed to
$\frac{\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\mathbf{R}\vec{\mathbf{p}}_{2})+(\mathbf{R}\vec{\mathbf{p}}_{1})^{T}(\mathbf{R}\vec{\mathbf{p}}_{2})\big{)}}{\int_{\mathbf{R}\vec{\mathbf{p}}_{3}\in\Omega}\exp\big{(}f^{Q}(\mathbf{R}\vec{\mathbf{p}}_{1})*f^{K}(\mathbf{R}\vec{\mathbf{p}}_{3})+(\mathbf{R}\vec{\mathbf{p}}_{1})^{T}(\mathbf{R}\vec{\mathbf{p}}_{3})\big{)}d\mathbf{R}\vec{\mathbf{p}}_{3}}.$ (37)
Due to the spherical symmetry, we can still eliminate the rotation $\mathbf{R}$ acting on $\mathbf{p}_{2}$ and $\mathbf{p}_{3}$. Therefore, after incorporating the Relative Orientation Embedding, EST retains its equivariance.
Appendix C Details of MoE
C.1 Mixture of Experts in Language Models
Recently, empirical evidence consistently demonstrates that increased model parameters and computational resources yield performance gains in language models when sufficient training data is available. However, scaling models to extreme sizes incurs prohibitive computational costs. The Mixture-of-Experts (MoE) architecture has emerged as a promising solution to this dilemma. By enabling parameter scaling while maintaining moderate computational requirements, MoE architectures have shown particular success when integrated with Transformer frameworks. These implementations have successfully scaled language models to substantial sizes while preserving performance advantages.
C.2 Mixture of Experts for Transformers
We begin with a standard Transformer language model architecture, which comprises $Y$ stacked Transformer blocks:
$\mathbf{u}^{y}$ $=\operatorname{Self-Att}\left(\mathbf{h}^{m-1}\right)+\mathbf{h}^{m-1},$ (38)
$\mathbf{h}^{y}$ $=\operatorname{FFN}\left(\mathbf{u}^{y}\right)+\mathbf{u}^{y},$ (39)
where $\operatorname{Self-Att}(\cdot)$ denotes the self-attention module, $\operatorname{FFN}(\cdot)$ denotes the Feed-Forward Network (FFN), $\mathbf{u}^{y}$ are the hidden states of all tokens after the $y$-th attention module, and $\mathbf{h}^{y}\in\mathbb{R}^{d}$ is the output hidden state after the $y$-th Transformer block. layer normalization is omitted for brevity. The MoE architecture substitutes FFN layers in Transformers with MoE layers and each MoE layer comprises multiple structurally identical experts. Each token is dynamically assigned to several experts based on learned routing probabilities: If the $y$-th FFN is substituted with an MoE layer, the computation for its output hidden state $\mathbf{h}^{y}$ can be expressed as:
$\mathbf{h}^{y}$ $=\sum_{i=1}^{N}\left({g_{i}\operatorname{FFN}_{i}\left(\mathbf{u}^{y}\right)}\right)+\mathbf{u}^{y},$ (40)
$g_{i}$ $=\begin{cases}s_{i},&s_{i}\in\operatorname{Topk}(\{s_{j}|1\leq j\leq N\},K),\\ 0,&\text{otherwise},\end{cases}$ (41)
$s_{i}$ $=\operatorname{Softmax}_{i}\left({(\mathbf{u}^{y})}^{T}\mathbf{e}_{i}^{y}\right),$ (42)
where $N$ denotes the total number of experts, $\operatorname{FFN}_{i}(\cdot)$ is the $i$-th expert FFN, $g_{i}$ denotes the gate value for the $i$-th expert, $s_{i}$ denotes the token-to-expert affinity, $\operatorname{Topk}(\cdot,K)$ denotes the set comprising $K$ highest affinity scores among those calculated for all $N$ experts, and $\mathbf{e}_{i}^{y}$ is the learnable parameters representing the centroid of the $i$-th expert in the $y$-th layer. The sparsity property ($K\ll N$) ensures computational efficiency by restricting each token to interact with only $K$ experts.
C.3 Mixture of Hybrid Experts in EST
Inspired by MoE of language models, we developed the mixture of hybrid experts for EST. Its computation can be expressed as:
$\tilde{\mathbf{m}}$ $=\sum_{i=1}^{N_{steerable}}\left({g_{i}\operatorname{SteerableFFN}_{i}\left(\mathbf{m}\right)}\right)+\sum_{j=1}^{N_{spherical}}\left({g_{j}\operatorname{SphericalFFN}_{i}\left(\mathbf{m}\right)}\right)+\mathbf{m},$ (43)
$g_{i}$ $=\begin{cases}s_{i},&s_{i}\in\operatorname{Topk}(\{s_{k}|1\leq k\leq N_{steerable}\},K),\\ 0,&\text{otherwise},\end{cases}$ (44)
$g_{j}$ $=\begin{cases}s_{j},&s_{j}\in\operatorname{Topk}(\{s_{k}|1\leq k\leq N_{spherical}\},K),\\ 0,&\text{otherwise},\end{cases}$ (45)
$s_{i},s_{j}$ $=\mathrm{split}\left(\operatorname{Softmax}\left({\mathbf{m}^{(0)}}\mathbf{W}\right)\right),$ (46)
where $\mathbf{m}$ and $\tilde{\mathbf{m}}$ denote the input and output message, respectively, $\mathbf{m}^{(0)}$ denotes the invariant part of message, $\mathbf{W}\in\mathbb{R}^{C\times(N_{steerable}+N_{spherical})}$ represents the learnable expert centroids. Here, we omit the symbol of layer order for simplicity.
C.4 Ablation Studies
In terms of MoE, there are three key hyper-parameters, including expert count, routing sparsity (top-K), and load balancing. Our ablation studies begin with the easiest steerable MoE using the QM9 dataset and Equiformer baseline.
  • Expert count: We assessed model performance across expert counts ranging from 1 to 64. Performance improved up to 10 experts, reaching optimal results at that point (Table 6). Meanwhile, excessive expert count leads to a significant increase in time cost while the improvement in performance may not necessarily continue. In contrast, setting the number of expert as 10 offers the right balance between performance improvement and time cost.
  • Routing Sparsity: We evaluated different levels of sparsity by varying the K values in the top-K router, while increased sparsity (lower K values) degraded performance (Table 7).
  • Load Balancing: The auxiliary balance loss was added to the training loss with varying ratios (0 to 0.1). However, introducing balance loss consistently reduced model performance (Table 8).
The discrepancy between our findings and typical language model observations regarding sparsity and balancing can be attributed to several factors. While sparsity and balancing mechanisms generally enhance efficiency in standard language models, the limited size of the QM9 dataset plays a crucial role in this divergence. Introducing more experts increases model complexity, which can lead to overfitting rather than improved generalization, particularly when training data is scarce. Furthermore, given our limited number of experts, the necessity for stringent sparsity enforcement and elaborate load balancing mechanisms is diminished. In such scenarios, employing a dense MoE configuration tends to yield optimal performance.
Table 6: Performance and time cost depending on expert count when predicting $\alpha$ on QM9.
Number of experts 1(w/o MoE) 2 4 8 10 16 32 64
Test MAE $\left(\mathrm{bohr}^{3}\right)$ 0.0466 0.0461 0.0503 0.0437 0.0424 0.0449 0.0438 0.0443
Time per epoch ($s$) 373.69 391.93 416.53 472.80 500.71 576.01 893.20 1664.93
Table 7: Effect of routing sparsity for predicting $\alpha$ on QM9.
K 2 3 5 10 (Dense)
Test MAE $\left(\mathrm{bohr}^{3}\right)$ 0.04338 0.04349 0.04472 0.0424
Table 8: Impact of balance loss ratio for predicting $\alpha$ on QM9.
Ratio 0 (w/o balance loss) 0.001 0.01 0.1
Test MAE $\left(\mathrm{bohr}^{3}\right)$ 0.0424 0.04369 0.04369 0.04361
Appendix D Details of Experiments and Supplementary Experiments
D.1 Implementation Details of Baselines
Table 9: Hyper-parameters for the EST model setting on OC20 S2EF and OC20 IS2RE experiments.
Hyper-parameters S2EF-ALL IS2RE
Optimizer AdamW AdamW
Learning rate scheduling Cosine learning rate with linear warmup Cosine learning rate with linear warmup
Warmup epochs $0.01$ 22
Maximum learning rate $4\times 10^{-4}$ $2\times 10^{-4}$
Batch size $256$ $32$
Number of epochs 33 2020
Weight decay $\times 10^{-3} 1×1$$\times 10^{-3}$
Dropout rate $0.1$ $0.2$
Energy coefficient $\lambda_{E}$ 44 11
Force coefficient $\lambda_{F}$ $100$ -
Gradient clipping norm threshold $100$ $100$
Model EMA decay $0.999$ $0.999$
Cutoff radius (Å) $12$ 55
Maximum number of neighbors 2020 $500$
Number of radial bases $600$ $128$
Dimension of hidden scalar features in radial functions $128$ $64$
Maximum degree $L_{max}$ 66 11
Maximum order $M_{max}$ 22 11
Number of Layers 88 66
Node embedding dimension $128$ $(256,l=0),(128,l=1)$
Intermediate dimension during the Fourier transform $128$ $256$
Intermediate dimension and the number of steerable FFN $128,6$ $[(768,l=0),(384,l=1)],5$
Intermediate dimension and the number of spherical FFN $512,4$ $512,5$
Number of spherical point samples $128$ $128$
In the S2EF experiment, the results of baselines in Table 1 follow [17], where each model is trained in official configuration. Most of these configurations can be found in Fairchem repository, and we also follow its code framework to construct our OC20 experiments. In the IS2RE experiment, the results in Table 2 of baselines follow [12] and [15]. In the QM9 experiment, the results in Table 3 of baselines follow [17] and [23].
D.2 Implementation Details of EST Experiments
S2EF and IS2RE
In our experiments on OC20, we adopt two hybrid models: S2EF combines the message module from EquiformerV2 with the update module of EST, while IS2RE combines the message module of Equiformer with the update module of EST. To ensure a fair comparison, all training configurations are aligned with those of the original EquiformerV2 and Equiformer. The hyperparameters specific to the EST architecture include the number of spherical sampling points, the number and dimension of experts in the steerable FFN, and the number of experts in the spherical FFN. Detailed configurations for all models are summarized in Table 9. The experiment on S2EF is conducted on 32 NVIDIA A100 GPUs and the experiment on S2EF is conducted on 8 NVIDIA A100 GPUs.
Table 10: Hyper-parameters for QM9 dataset.
Hyper-parameters EST EST (with GA)
Optimizer AdamW AdamW
Learning rate scheduling Cosine learning rate with Cosine learning rate with
linear warmup linear warmup
Warmup steps 55 55
Maximum learning rate $5\times 10^{-4}$, $2\times 10^{-4}$ $5\times 10^{-4}$, $1.5\times 10^{-4}$
Batch size $128,64$ $128,64$
Max training epochs $350,700$ $300,600$
Weight decay $5\times 10^{-3},0$ $5\times 10^{-3},0$
Dropout rate $0.0,0.2$ $0.0,0.2$
Number of radial bases $128$ for Gaussian radial basis $8$ for radial bessel basis
Cutoff radius (Å) 55 55
$L_{max}$ 22 22
Number of layers 66 66
Node dimension $(128,l=0),(64,l=1),(32,l=2)$ $128$
Spherical harmonics embedding dimension $(1,l=0),(1,l=1),(1,l=2)$ $(1,l=0),(1,l=1),(1,l=2)$
Intermediate dimension during the $128$ $128$
Fourier transform
Intermediate dimension and the number of $5,128$ $6,(768,l=0),(384,l=1)$
steerable FFN
Intermediate dimension and the number of $5,512$ $4,512$
spherical FFN
Number of spherical point samples $200,128$ $128$
QM9
In the QM9 experiments, we design two EST variants: 1) a fully EST-based architecture with both message and update blocks; 2) a hybrid model combining the message block of Equiformer and the update block of EST . To ensure a fair comparison with state-of-the-art methods (Equiformer and EquiformerV2 ), we adopt similar configurations as shown in Table 10. Note that Equiformer employs two different configurations for the following properties: $\alpha,\Delta\varepsilon,\varepsilon_{\text{HOMO}},\varepsilon_{\text{LUMO}},\mu,C_{v},R^{2},\text{ZPVE}$, and $G,H,U,U_{0}$. We follow the same strategy in our experiments. Comparisons with another SOTA model, GotenNet [23], are conducted under different configurations, and the details are provided in Section D.3.
D.3 Comparison between GotenNet and EST
Comparing with GotenNet, a variant of EST was designed incorporating the GATA message block and substituting the original updating block with our proposed mixture of hybrid experts. The steerable part of the hybrid expert maintain the architecture of EQFF in GotenNet, while the EST updating module serves as the spherical part used in parallel. To maintain a lightweight design, we simply integrate one steerable expert with one spherical expert. Our model adopts the same number of layers as GotenNet but achieves better performance across 6 tasks (table 11).
To ensure a fair comparison, we use training configurations similar to those of the original GotenNet (see Table 12), including a batch size of $32$ and a training schedule of $1000$ epochs. It is notable that several training hyper-parameters of EST differ from those used of GotenNet. For example, GotenNet specifies 10,000 warm-up steps, whereas our model sets this to 0. Additionally, we employed MAE as the loss function, while MSE was used for GotenNet instead. Properties where EST performs poorly may stem from such subtle differences. For example, EST outperforms GotenNet on both $\varepsilon_{HOMO}$ and $\varepsilon_{LUMO}$, but it underperforms significantly on the $\Delta\varepsilon$ property, which is the gap between $\varepsilon_{HOMO}$ and $\varepsilon_{LUMO}$. We suspect that this might be due to the absence of a warmup stage, causing the model to converge to a suboptimal local minimum.
Table 11: Results on QM9 dataset for various properties.
Task $\alpha$ $\Delta\varepsilon$ $\varepsilon_{HOMO}$ $\varepsilon_{LUMO}$ $\mu$ $C_{v}$ $G$ $H$ $R^{2}$ $U$ $U_{0}$ ZPVE
Units $bohr^{3}$ meV meV meV D cal/(mol K) meV meV $bohr^{3}$ meV meV meV
GotenNet (4 layers) .033 21.2 16.9 13.9 .0075 .020 5.50 3.70 0.029 3.67 3.71 1.09
EST (4 layers with GATA) .030 31.2 16.7 13.3 .0070 .019 5.31 4.02 0.029 3.96 3.94 1.16
Table 12: Hyper-parameters for the EST and GotenNet models on QM9 experiments. Red highlights the configurations that differ.
Hyper-parameters EST (with GATA) GotenNet
Loss function MAE MSE
Learning rate scheduling reduce on plateau linear warmup with reduce on plateau
Warmup steps 0 $10000$
Weight decay 0 $0.01$
Optimizer AdamW AdamW
Maximum learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Batch size $32$ $32$
Max training epochs $1000$ $1000$
Dropout rate $0.1$ $0.1$
Number of RBFs $64$ $64$
Cutoff radius (Å) 55 55
$L_{max}$ 22 22
Number of layers 44 44
Node dimension $256$ $256$
Edge dimension $256$ $256$
Number of attention heads 88 88
Intermediate dimension and the number of $512,1$ $512,1$
steerable FFN
Intermediate dimension and the number of $512,1$ -
spherical FFN
Number of spherical point samples $64$ -
D.4 Supplementary Experiments
Building Blocks
We conducted ablation studies to validate the effectiveness of individual building blocks within EST. Specically, in Table 13, removing the steerable FFN means all experts are replaced by the spherical FFN, removing the spherical FFN means all experts are replaced by the steerable FFN and removing FL Sampling refers to using the Fourier transform from e3nn instead. We used the prediction of the $\alpha$ property on QM9 as the core task for evaluation. As shown in Table 13, all components positively contribute to the overall performance. We draw several conclusions:
  1. 1.
    The SA module with layer normalization effectively improves overall performance.
  2. 2.
    Mixing steerable and spherical experts helps strike a balance between equivariance and expressive power, leading to better generalization performance.
  3. 3.
    FL Sampling is crucial for EST. Disrupting equivariance without it significantly harms the results on QM9.
Table 13: Ablation studies for modules in HDGNN.
Building blocks in EST $\alpha$ MAE
LayerNorm SA Steerable FFN Spherical FFN FL Sampling $bohr^{3}$
- 0.046
- - 0.044
- 0.045
- - - 0.042
- 0.053
0.041